OpenStudy (kjay):

Assume that y varies inversely as x. If x = 12 when y = 9, what is x when y = -3?

OpenStudy (lgbasallote):

do you have an answer?

OpenStudy (anonymous):

y varies in inversely means y is the denominator like: \[\large x = k \times \frac{1}{y} \implies x = \frac{k}{y}\] Now x = 12 and y = 9 Plug in the values and find k first..

OpenStudy (lgbasallote):

you can also do \[y = \frac kx\] hehe

OpenStudy (anonymous):

Ha ha ha..

OpenStudy (anonymous):

After finding k, use the same formula to find x and put y = -3 and k = as you got..